# Small-time asymptotics and the emergence of complex singularities for the KdV equation

### v1

@@ Abstract
Keywords: KdV equation, complex-plane singularities, exponential asymptotics, matched asymptotic expansions, Stokes phenomenon, dispersive waves, solitary waves, Painlevé II, tritronquée solutions, transseries.
@@ 1 Introduction
$u_{t}+6uu_{x}+u_{xxx}=0,\quad u(x,0)=u_{0}(x),\quad x\in\mathbb{R},$ (1)
is to treat the analytic continuation of real solutions of (1) to the complex-$x$ plane, in part because the complex singularities of the solution are intimately related to the propagation of real-valued dispersive waves and solitons. In this spirit, Kruskal [56], Thickstun [79] and Bona & Weissler [11] study the role of complex-plane singularities of solutions in soliton interactions, while Airault et al. [4] and Deconinck & Segur [29] conduct related studies for elliptic solutions of the KdV equation. Other more recent research focuses on applying numerical methods to track singularities of solutions of the KdV equation, especially for periodic initial conditions or small dispersion [13, 15, 44, 85].
Following these studies, we are also interested in the evolution of complex-variable singularities of (the analytic continuation of) solutions to (1), in our case with a focus on the limit $t\rightarrow 0^{+}$. The small-time limit is important because singularities cannot be born at finite times; therefore, the initial number, type and motion of singularities can provide a solid indication of their behaviour for later times. Apart from being of interest in their own right, our study of singularity behaviour in the complex plane, especially those singularities closest to the real axis, has facilitated new results for the real-valued problem, including a short-time asymptotic description of the dispersive waves.
@@
Returning to our KdV problem, we shall be studying the analytic continuation of solutions of (1). These complex-valued solutions will have singularities in the complex-$x$ plane that are all double poles with principal part 2-2. For each such complex singularity $x=s(t)$, the local behaviour must be
$u\sim-\frac{2}{(x-s(t))^{2}}\quad\mbox{as}\quad x\rightarrow s(t),$ (2)
@@
$u_{0}\sim\frac{A_{0}}{(x-x_{0})^{2}}\quad\mbox{as}\quad x\rightarrow x_{0},$ (3)
where $A_{0}$ may be complex (although in this paper the examples we include have real values of $A_{0}$). Here, each $x_{0}\notin\mathbb{R}$ will be one of a complex-conjugate pair, since we want $u_{0}$ to be real for $x\in\mathbb{R}$. We shall address the broad question of how solutions with double poles like (2) emerge from initial conditions with (3), especially in terms of the small-time dynamics. The analytical tools we employ are based on formal asymptotics, including exponential asymptotics (using the approach in [16, 70]) and matched asymptotic expansions in the limit $t\rightarrow 0^{+}$ (cf. [28, 37, 82] and others), analyses of the Stokes phenomenon and transseries to locate poles of the inner problem near $x_{0}$ (like [24, 60]), while the computational tools include direct numerical solution of the real problem, and a pole-solver algorithm to analyse the inner problem numerically [39, 40]. Note that analytic initial conditions for (1) guarantee solutions that are analytic for $t>0$ [46, 49].
One reason for starting with analytic initial conditions with (3) is that it turns out the matched asymptotic expansions for the limit $t\rightarrow 0^{+}$ near each $x_{0}$ are much easier to deal with than if the singularity of $u_{0}$ is a different type (i.e., other than a double pole). Another reason for focussing on initial conditions with (3) is to observe the difference between having initial double poles with $A_{0}=-2$ and the more general case $A_{0}\neq-2$. Finally, the very well-studied $\mathrm{sech}^{2}$-type initial conditions have double poles as in (3), namely
$u_{0}=-A_{0}\,\mathrm{sech}^{2}x,\quad x_{0}=\left(n+{\textstyle\frac{1}{2}}\right)\pi\mathrm{i},\quad A_{0}\in\mathbb{R},\quad n\in\mathbb{Z},$ (4)
providing further motivation for (3). Properties of solutions of (1) with (4) are well known, mostly via the application of inverse-scattering techniques [3, 34]. For $A_{0}<0$, the profile $u_{0}$ in (4) on the real line is a “hump” centred at $x=0$, and the subsequent solution of (1) will in general involve a finite number of solitons propagating to the right, together with dispersive radiation characterised by waves propagating very quickly to the left, as illustrated in figure 1. For the special cases $A_{0}=-N(N+1)$, $N\in\mathbb{N}$, there are exact $N$-soliton solutions of (1) with (4), which will be of interest as test cases. The 1-soliton solution, $N=1$ ($A_{0}=-2$), is a exceptional case that falls outside our analysis for $A_{0}\neq-2$, as we shall explain. On the other hand, for $A_{0}>0$ the shape is a “dip”, with solutions of (1) involving dispersive waves only. One of the goals of our study is to attempt to identify links between complex-plane singularities of solutions and the qualitative behaviour of the solutions on the real line.
@@
Figure 1: Numerical solutions of the KdV model (1) on the real line using the $\mathrm{sech}^{2}$-type initial condition (4), showing dispersive waves propagating to the left. Calculations are performed using the spin command [66] in Chebfun [35]. For (a),(b) since $0<-A_{0}<2$, there are no solitons moving to the right; (c) since $2<-A_{0}<6$, there is one soliton; (d) since $6<-A_{0}<12$, there are two solitons. Note, although hard to see on this scale, the dispersive waves are roughly the same size for each of these four examples, while the height of the initial hump $u_{0}(0)=-A_{0}$ of course increases as $A_{0}$ increases.
@@
The outline of our paper is as follows. Various aspects of the small-time analysis for the generic case (1) with (3) and $A_{0}\neq-2$ are provided in section 2. The starting point in subsection 2.1 is an “outer” expansion in powers of $t$, which is not able to describe the dispersive waves on the real line, as their amplitudes are exponentially small in $t$ and hence are formally smaller than each term $t^{n}u_{n}(x)$ in the algebraic series. This is reflected in the outer expansion being divergent, with the leading-order term $u_{0}(x)$ having singularities at $x=x_{0}$ off the real axis (of the form (3)). Thus, we require techniques in exponential asymptotics and the Stokes phenomenon to extract details of the dispersive waves in the small-time limit, with our analysis in subsection 2.2 predicting that the dispersive waves behave, to leading order, as
$U_{\mbox{\footnotesize dis}}\sim-\frac{2}{3^{3/4}\pi^{1/2}}\cos\left(\frac{\pi}{2}\sqrt{1-4A_{0}}\right)\frac{(-x)^{1/4}\,\mathrm{e}^{-y_{0}(-x/3t)^{1/2}}}{t^{3/4}}\,\cos\left(\frac{2(-x)^{3/2}}{3(3t)^{1/2}}-\frac{\pi}{4}\right)$ (5)
@@
as $x\rightarrow$-$\infty$, $t\rightarrow 0^{+}$, where $y_{0}=\mathrm{Im}(x_{0})$ (further details are provided in appendix A). It is noteworthy that the cosine term out the front of the expression in (5) has an explicit dependence on $A_{0}$, linking the amplitude of the dispersive waves to the strength of the double pole at $x=x_{0}$, with this term vanishing for the special cases $A_{0}=-N(N+1)$, where $N$ is a natural number. We test our exponential asymptotics in subsection 2.3 by comparing with numerical solutions of (1) for the initial conditions (4) and
$u_{0}=-\frac{4A_{0}}{(1+x^{2})^{2}},\quad x_{0}=\pm\mathrm{i},\quad A_{0}\in\mathbb{R},$ (6)
@@
(6) being worth including in part because it is not of the special class (4) and so therefore does not give rise to known exact solutions. Further, (6) is instructive because this initial condition has only one singularity, $x_{0}=\mathrm{i}$, in the upper half plane, so we avoid any possible complications from singularities born at multiple points in each half plane. An additional, carefully constructed, initial condition with a vanishing residue (namely (22)) is also used as a comparison.
@@
We continue our study of the small-time limit in section 3, where we consider an “inner” problem near $x=x_{0}$ in subsection 3.1. We show how the initial dynamics of the singularities of our KdV problem are governed by the inhomogeneous Painlevé II (PII) equation (subsection 3.2) with decreasing tritronquée solutions (subsection 3.3). In general, there are infinitely many singularities that emerge from each $x=x_{0}$; exceptional cases arise for $A_{0}=-N(N+1)$, whereby well-known rational solutions of PII (subsection 3.4) correspond to a finite number of singularities for KdV. (A summary of the effects of higher-order corrections to (3) is provided in appendix B, including some special cases.) To illustrate the singularity structure of our PII solutions, some numerical results are presented in subsection 3.5. To close section 3, we use a standard transseries approach in subsection 3.6 to derive approximations to the locations of the most important of these singularities.
@@
As concrete examples of KdV solutions, the widely-studied initial conditions (4) are used in section 4 to illustrate some of the global features of our small-time analysis from section 2, focussing on the role of rational solutions of PII in the small-time dynamics of $N$-soliton solutions. It is worth emphasising that, while the integrability of the KdV equation makes some of the calculations in various parts of our study more analytically amenable, our methodology should be broadly applicable to other dispersive wave equations. We allude to these observations in section 5, where we also summarise our findings more generally and present a discussion about a significant number of unresolved issues and open problems.
2 Small-time analysis for (1) with (3)
@@
As mentioned in Section 1, one of our main motivations is to understand the role of complex-plane singularities of KdV solutions, which for $t>0$ must be double poles of the form (2). In this section and in section 3, we concentrate on analytic initial conditions that also have double poles, as in (3), but with a different principal part to (2), namely $A_{0}\neq-2$. The methodology we employ is based in part on matched asymptotic expansions in the limit $t\rightarrow 0^{+}$, including an outer region away from complex singularities of the initial condition $u_{0}(x)$ and inner regions near these singularities, as well as exponential asymptotics that involves analysis from both the outer and inner regions. The present section summarises the exponential asymptotics, while a more detailed study of the inner regions is deferred until section 3.
@@ 2.1 Outer expansion away from $x=x_{0}$
Consider the KdV equation (1) together with an initial condition $u_{0}$ with the property (3), where $x_{0}$ (with $\mathrm{Im}(x_{0})=y_{0}>0$) is one of a complex-conjugate pair. Assuming that $u_{0}$ is real for $x\in\mathbb{R}$, then a consequence of this problem formulation is that we need consider the solution only on the real line and the upper half plane, with the understanding that the behaviour of the singularities in the lower half plane will be an appropriate reflection about the real-$x$ axis.
@@
$u\sim u_{0}(x)+tu_{1}(x)\quad\mbox{as}\quad t\rightarrow 0^{+}.$ (7)
By substituting into (1), we find
$u_{1}=-6u_{0}u_{0}^{\prime}-u_{0}^{\prime\prime\prime}.$ (8)
Therefore, we have the local behaviour
$u_{0}\sim\frac{A_{0}}{(x-x_{0})^{2}},\quad u_{1}\sim\frac{12A_{0}(A_{0}+2)}{(x-x_{0})^{5}}\quad\mbox{as}\quad x\rightarrow x_{0},$ (9)
noting that $u_{1}$ is three orders more singular than $u_{0}$ at $x=x_{0}$ (due to the third derivative in (8)). Thus, we see immediately that the outer expansion (7) applies along the real axis and in parts of the upper half complex plane away from $x=x_{0}$, but breaks down in regions where $u_{0}=\mathcal{O}(tu_{1})$, or, in other words, where
@@
This reasoning suggests there will be an inner problem near $x=x_{0}$, which we consider in subsection 3.1. (Note that the power-series expansion will also break down in a sector in the upper half plane bounded by anti-Stokes lines, as discussed below in subsections 2.2 and 3.3.)
@@ 2.2 Exponential asymptotics argument for limit $t\rightarrow 0^{+}$
The terms $u_{0}$ and $u_{1}$ in (7) are the first two in a divergent asymptotic expansion of the form
$u\sim\sum_{n=0}^{\infty}t^{n}u_{n}(x)\quad\mbox{as}\quad t\rightarrow 0^{+},$ (10)
whose divergence is caused by the singularities of the leading-order term $u_{0}(x)$ in the complex-$x$ plane. As such, the $u_{n}$ will be of the familiar factorial-over-power form for large $n$, due to repeatedly differentiating (three times) to obtain the next-order term. Thus, we expect there to be an exponential term in our asymptotic representation for $u$ in the limit $t\rightarrow 0^{+}$ that appears “beyond all orders” of the original power series (10) [16, 70]. This term will “switch on” across Stokes lines in the $x$ plane and, in particular, will affect the solution on the real line by providing an exponentially small correction term to (10), which we call $U_{\mbox{\footnotesize dis}}$.
Importantly, the emergence of dispersive waves that travel in the negative-$x$ direction can never be described by a power series (10), as the amplitude of these waves turns out to be exponentially small in time compared to each term $t^{n}u_{n}$. Thus, in order to approximate the dispersive wavetrain in the limit $t\rightarrow 0^{+}$, we must consider this exponential contribution and observe where it switches on across the real-$x$ axis. To demonstrate how this works, we follow the framework in [16], summarise the main results here, and provide further details in appendix A (see [17, 57, 58, 59, 80], for example, for other studies of wave motion using this framework of exponential asymptotics).
A crucial step is to analyse the late terms in (10), which satisfy
$(n+1)u_{n+1}=-6\sum_{j=0}^{n}u_{j}u^{\prime}_{n-j}-u_{n}^{\prime\prime\prime}.$ (11)
Due to the factorial over power divergence of (10), following Dingle [32] we apply the ansatz
$u_{n}\sim\frac{\mathcal{A}(x)\Gamma(2n+\gamma)}{\chi(x)^{2n+\gamma}}\quad\mbox{as}\quad n\rightarrow\infty,$ (12)
which, after substituting into (11), leads to
$\frac{1}{2}\chi=\chi^{\prime 3},\quad-\frac{\gamma}{2}\mathcal{A}=-(3\mathcal{A}\chi^{\prime}\chi^{\prime\prime}+3\mathcal{A}^{\prime}\chi^{\prime 2}).$ (13)
Here $\chi$ is the so-called singulant, which must vanish at singularities of $u_{0}$, and it suffices to consider the location $x=x_{0}$ (the full expansion of $u_{n}$ for large $n$ comprises a sum of terms of the form (12) associated with each singularity $x_{0}$). Thus, given $x_{0}$ is assumed to lie in the upper half plane, we explain in appendix A.1 that the appropriate solutions are
$\chi=-\frac{2}{3^{3/2}}(x-x_{0})^{3/2},\quad\mathcal{A}=\Lambda(x-x_{0})^{(\gamma-1)/2},$ (14)
@@
$u_{n}\sim\frac{\Lambda(-3^{3/2}/2)^{2n+\gamma}\Gamma(2n+\gamma)}{(x-x_{0})^{3n+\gamma+1/2}}\quad\mbox{as}\quad n\rightarrow\infty.$ (15)
For the initial conditions we are concerned with in this study, we have (3). To be consistent with this local behaviour near $x=x_{0}$, we must choose $\gamma=3/2$.
@@
$\pi\mathrm{i}\,\mathcal{A}\,t^{-\gamma/2}\mathrm{e}^{-\chi/t^{1/2}}=\pi\mathrm{i}\,\Lambda(x-x_{0})^{1/4}t^{-3/4}\mathrm{e}^{2(x-x_{0})^{3/2}/3(3t)^{1/2}}$ (16)
switches on as we cross the Stokes line $\mathrm{arg}(x-x_{0})=-2\pi/3$ from right to left (given that $x_{0}$ is in the upper half $x$ plane). This Stokes line is found by setting the singulant term $\chi$ to be real and positive. There will be another term like (16) switched on across a Stokes line born in the lower half plane, and together the sum of these provides the asymptotic behaviour of the dispersive waves on the real line in the small-time limit.
To determine the constant $\Lambda$ in (16), we need to match into an inner region near $x=x_{0}$, which we study below in subsection 3.1. The details for this matching are provided in appendix A.3, with the key result that
$\Lambda=\frac{\mathrm{i}}{3^{3/4}\pi^{3/2}}\cos\left(\frac{\pi}{2}\sqrt{1-4A_{0}}\right).$ (17)
A striking property of $\Lambda$ is that it vanishes for $A_{0}=-N(N+1)$, where $N\in\mathbb{N}$. Therefore, this leading-order result predicts that a train of dispersive waves is associated with each complex conjugate pair of singularities of the initial condition $u_{0}$, with the amplitude vanishing for $A_{0}=-N(N+1)$. This prediction is consistent with the $\mathrm{sech}^{2}$-type initial condition (4), for which it is known that there are no dispersive waves when $A_{0}$ takes on these triangular values (these are the $N$-soliton solutions, having flat tails). (More generally, if $A_{0}=-N(N+1)$ for an initial condition that is not of the special form (4), then parts of the required asymptotics to describe the dispersive waves will be different; we include a discussion on such cases in appendix A.4.)
To take an example, consider an initial condition that has a singularity lying on the imaginary axis, which we write as $x_{0}=\mathrm{i}y_{0}$. Our analysis predicts that the term (16) switches on across the Stokes line $\mathrm{arg}(x-\mathrm{i}y_{0})=-2\pi/3$ in the upper half $x$ plane. We would also observe an analogous term that switches on across $\mathrm{arg}(x+\mathrm{i}y_{0})=2\pi/3$ in the lower half plane. These Stokes lines both hit the real axis at $x=-y_{0}/\sqrt{3}$. Thus, putting it together, we find the exponential contribution to our asymptotic solution on the real-$x$ axis is of the form

### v2

@@ Abstract
keywords:
KdV equation, complex-plane singularities, exponential asymptotics, matched asymptotic expansions, Stokes phenomenon, dispersive waves, solitary waves, Painlevé II, tritronquée solutions, transseries
journal: Journal of Nonlinear Wavesarticletype: RESEARCH ARTICLE
[…]
2026
{Frontmatter}
[…]
Scott W. McCue et al.
@@ 1 Introduction
$u_{t}+6uu_{x}+u_{xxx}=0,\quad u(x,0)=u_{0}(x),\quad x\in\mathbb{R},$ (1.1)
is to treat the analytic continuation of real solutions of (1.1) to the complex-$x$ plane, in part because the complex singularities of the solution are intimately related to the propagation of real-valued dispersive waves and solitons. In this spirit, Kruskal [65], Thickstun [90] and Bona & Weissler [13] study the role of complex-plane singularities of solutions in soliton interactions, while Airault et al. [5] and Deconinck & Segur [34] conduct related studies for elliptic solutions of the KdV equation. Other more recent research focuses on applying numerical methods to track singularities of solutions of the KdV equation, especially for periodic initial conditions or small dispersion [15, 18, 50, 96].
Following these studies, we are also interested in the evolution of complex-variable singularities of (the analytic continuation of) solutions to (1.1), in our case with a focus on the limit $t\rightarrow 0^{+}$. The small-time limit is important because singularities cannot be born at finite times; therefore, the initial number, type and motion of singularities can provide a solid indication of their behaviour for later times. Apart from being of interest in their own right, our study of singularity behaviour in the complex plane, especially those singularities closest to the real axis, has facilitated new results for the real-valued problem, including a short-time asymptotic description of the dispersive waves.
@@
Returning to our KdV problem, we shall be studying the analytic continuation of solutions of (1.1). These complex-valued solutions will have singularities in the complex-$x$ plane that are all double poles with principal part 2-2. For each such complex singularity $x=s(t)$, the local behaviour must be
$u\sim-\frac{2}{(x-s(t))^{2}}\quad\mbox{as}\quad x\rightarrow s(t),$ (1.2)
@@
$u_{0}\sim\frac{A_{0}}{(x-x_{0})^{2}}\quad\mbox{as}\quad x\rightarrow x_{0},$ (1.3)
where $A_{0}$ may be complex (although in this paper the examples we include have real values of $A_{0}$). Here, each $x_{0}\notin\mathbb{R}$ will be one of a complex-conjugate pair, since we want $u_{0}$ to be real for $x\in\mathbb{R}$. We shall address the broad question of how solutions with double poles like (1.2) emerge from initial conditions with (1.3), especially in terms of the small-time dynamics. The analytical tools we employ are based on formal asymptotics, including exponential asymptotics (using the approach in [20, 79]) and matched asymptotic expansions in the limit $t\rightarrow 0^{+}$ (cf. [33, 42, 93] and others), analyses of the Stokes phenomenon and transseries to locate poles of the inner problem near $x_{0}$ (like [29, 66]), while the computational tools include direct numerical solution of the real problem, and a pole-solver algorithm to analyse the inner problem numerically [45, 46]. Note that analytic initial conditions for (1.1) guarantee solutions that are analytic for $t>0$ [52, 55].
One reason for starting with analytic initial conditions with (1.3) is that it turns out the matched asymptotic expansions for the limit $t\rightarrow 0^{+}$ near each $x_{0}$ are much easier to deal with than if the singularity of $u_{0}$ is a different type (i.e., other than a double pole). Another reason for focussing on initial conditions with (1.3) is to observe the difference between having initial double poles with $A_{0}=-2$ and the more general case $A_{0}\neq-2$. Finally, the very well-studied $\mathrm{sech}^{2}$-type initial conditions have double poles as in (1.3), namely
$u_{0}=-A_{0}\,\mathrm{sech}^{2}x,\quad x_{0}=\left(n+{\textstyle\frac{1}{2}}\right)\pi\mathrm{i},\quad A_{0}\in\mathbb{R},\quad n\in\mathbb{Z},$ (1.4)
providing further motivation for (1.3). Properties of solutions of (1.1) with (1.4) are well known, mostly via the application of inverse-scattering techniques [1, 39]. For $A_{0}<0$, the profile $u_{0}$ in (1.4) on the real line is a “hump” centred at $x=0$, and the subsequent solution of (1.1) will in general involve a finite number of solitons propagating to the right, together with dispersive radiation characterised by waves propagating very quickly to the left, as illustrated in figure 1. For the special cases $A_{0}=-N(N+1)$, $N\in\mathbb{N}$, there are exact $N$-soliton solutions of (1.1) with (1.4), which will be of interest as test cases. The 1-soliton solution, $N=1$ ($A_{0}=-2$), is an exceptional case that falls outside our analysis for $A_{0}\neq-2$, as we shall explain. On the other hand, for $A_{0}>0$ the shape is a “dip”, with solutions of (1.1) involving dispersive waves only. One of the goals of our study is to attempt to identify links between complex-plane singularities of solutions and the qualitative behaviour of the solutions on the real line.
@@
Figure 1: Numerical solutions of the KdV model (1.1) on the real line using the $\mathrm{sech}^{2}$-type initial condition (1.4), showing dispersive waves propagating to the left. Calculations are performed using the spin command [75] in Chebfun [40]. For (a),(b) since $0<-A_{0}<2$, there are no solitons moving to the right; (c) since $2<-A_{0}<6$, there is one soliton; (d) since $6<-A_{0}<12$, there are two solitons. Note, although hard to see on this scale, the dispersive waves are roughly the same size for each of these four examples, while the height of the initial hump $u_{0}(0)=-A_{0}$ of course increases as $A_{0}$ increases.
@@
The outline of our paper is as follows. Various aspects of the small-time analysis for the generic case (1.1) with (1.3) and $A_{0}\neq-2$ are provided in section 2. The starting point in subsection 2.1 is an “outer” expansion in powers of $t$, which is not able to describe the dispersive waves on the real line, as their amplitudes are exponentially small in $t$ and hence are formally smaller than each term $t^{n}u_{n}(x)$ in the algebraic series. This is reflected in the outer expansion being divergent, with the leading-order term $u_{0}(x)$ having singularities at $x=x_{0}$ off the real axis (of the form (1.3)). Thus, we require techniques in exponential asymptotics and the Stokes phenomenon to extract details of the dispersive waves in the small-time limit, with our analysis in subsection 2.2 predicting that the dispersive waves behave, to leading order, as
$U_{\mbox{\footnotesize dis}}\sim-\frac{2}{3^{3/4}\pi^{1/2}}\cos\left(\frac{\pi}{2}\sqrt{1-4A_{0}}\right)\frac{(-x)^{1/4}\,\mathrm{e}^{-y_{0}(-x/3t)^{1/2}}}{t^{3/4}}\,\cos\left(\frac{2(-x)^{3/2}}{3(3t)^{1/2}}-\frac{\pi}{4}\right)$ (1.5)
@@
as $x\rightarrow$-$\infty$, $t\rightarrow 0^{+}$, where $y_{0}=\mathrm{Im}(x_{0})$ (further details are provided in appendix A). It is noteworthy that the cosine term out the front of the expression in (1.5) has an explicit dependence on $A_{0}$, linking the amplitude of the dispersive waves to the strength of the double pole at $x=x_{0}$, with this term vanishing for the special cases $A_{0}=-N(N+1)$, where $N$ is a natural number. We test our exponential asymptotics in subsection 2.3 by comparing with numerical solutions of (1.1) for the initial conditions (1.4) and
$u_{0}=-\frac{4A_{0}}{(1+x^{2})^{2}},\quad x_{0}=\pm\mathrm{i},\quad A_{0}\in\mathbb{R},$ (1.6)
@@
(1.6) being worth including in part because it is not of the special class (1.4) and so therefore does not give rise to known exact solutions. Further, (1.6) is instructive because this initial condition has only one singularity, $x_{0}=\mathrm{i}$, in the upper half plane, so we avoid any possible complications from singularities born at multiple points in each half plane. An additional, carefully constructed, initial condition with a vanishing residue (namely (2.16)) is also used as a comparison.
@@
We continue our study of the small-time limit in section 3, where we consider an “inner” problem near $x=x_{0}$ in subsection 3.1. We show how the initial dynamics of the singularities of our KdV problem are governed by the inhomogeneous Painlevé II (PII) equation (subsection 3.2) with decreasing tritronquée solutions (subsection 3.4). In general, there are infinitely many singularities that emerge from each $x=x_{0}$; exceptional cases arise for $A_{0}=-N(N+1)$, whereby well-known rational solutions of PII (subsection 3.5) correspond to a finite number of singularities for KdV. (A summary of the effects of higher-order corrections to (1.3) is provided in appendix B, including some special cases.) To illustrate the singularity structure of our PII solutions, some numerical results are presented in subsection 3.6. To close section 3, we use a standard transseries approach in subsection 3.7 to derive approximations to the locations of the most important of these singularities.
@@
As concrete examples of KdV solutions, the widely-studied initial conditions (1.4) are used in section 4 to illustrate some of the global features of our small-time analysis from section 2, focussing on the role of rational solutions of PII in the small-time dynamics of $N$-soliton solutions. It is worth emphasising that, while the integrability of the KdV equation makes some of the calculations in various parts of our study more analytically amenable, our methodology should be broadly applicable to other dispersive wave equations. We allude to these observations in section 5, where we also summarise our findings more generally and present a discussion about a significant number of unresolved issues and open problems.
2 Small-time analysis for (1.1) with (1.3)
@@
As mentioned in Section 1, one of our main motivations is to understand the role of complex-plane singularities of KdV solutions, which for $t>0$ must be double poles of the form (1.2). In this section and in section 3, we concentrate on analytic initial conditions that also have double poles, as in (1.3), but with a different principal part to (1.2), namely $A_{0}\neq-2$. The methodology we employ is based in part on matched asymptotic expansions in the limit $t\rightarrow 0^{+}$, including an outer region away from complex singularities of the initial condition $u_{0}(x)$ and inner regions near these singularities, as well as exponential asymptotics that involves analysis from both the outer and inner regions. The present section summarises the exponential asymptotics, while a more detailed study of the inner regions is deferred until section 3.
@@ 2.1 Outer expansion away from $x=x_{0}$
Consider the KdV equation (1.1) together with an initial condition $u_{0}$ with the property (1.3), where $x_{0}$ (with $\mathrm{Im}(x_{0})=y_{0}>0$) is one of a complex-conjugate pair. Assuming that $u_{0}$ is real for $x\in\mathbb{R}$, then a consequence of this problem formulation is that we need consider the solution only on the real line and the upper half plane, with the understanding that the behaviour of the singularities in the lower half plane will be an appropriate reflection about the real-$x$ axis.
@@
$u\sim u_{0}(x)+tu_{1}(x)\quad\mbox{as}\quad t\rightarrow 0^{+}.$ (2.1)
By substituting into (1.1), we find
$u_{1}=-6u_{0}u_{0}^{\prime}-u_{0}^{\prime\prime\prime},$ (2.2)
where the primes (here and throughout the document) denote differentiation with respect to $x$. Therefore, we have the local behaviour
$u_{0}\sim\frac{A_{0}}{(x-x_{0})^{2}},\quad u_{1}\sim\frac{12A_{0}(A_{0}+2)}{(x-x_{0})^{5}}\quad\mbox{as}\quad x\rightarrow x_{0},$ (2.3)
noting that $u_{1}$ is three orders more singular than $u_{0}$ at $x=x_{0}$ (due to the third derivative in (2.2)). Thus, we see immediately that the outer expansion (2.1) applies along the real axis and in parts of the upper half complex plane away from $x=x_{0}$, but breaks down in regions where $u_{0}=\mathcal{O}(tu_{1})$, or, in other words, where
@@
This reasoning suggests there will be an inner problem near $x=x_{0}$, which we consider in subsection 3.1. (Note that the power-series expansion will also break down in a sector in the upper half plane bounded by anti-Stokes lines, as discussed below in subsections 2.2 and 3.4.)
@@ 2.2 Exponential asymptotics argument for limit $t\rightarrow 0^{+}$
The terms $u_{0}$ and $u_{1}$ in (2.1) are the first two in a divergent asymptotic expansion of the form
$u\sim\sum_{n=0}^{\infty}t^{n}u_{n}(x)\quad\mbox{as}\quad t\rightarrow 0^{+},$ (2.4)
whose divergence is caused by the singularities of the leading-order term $u_{0}(x)$ in the complex-$x$ plane. As such, the $u_{n}$ will be of the familiar factorial-over-power form for large $n$, due to repeatedly differentiating (three times) to obtain the next-order term. Thus, we expect there to be an exponential term in our asymptotic representation for $u$ in the limit $t\rightarrow 0^{+}$ that appears “beyond all orders” of the original power series (2.4) [20, 79]. This term will “switch on” across Stokes lines in the $x$ plane and, in particular, will affect the solution on the real line by providing an exponentially small correction term to (2.4), which we call $U_{\mbox{\footnotesize dis}}$.
Importantly, the emergence of dispersive waves that travel in the negative-$x$ direction can never be described by a power series (2.4), as the amplitude of these waves turns out to be exponentially small in time compared to each term $t^{n}u_{n}$. Thus, in order to approximate the dispersive wavetrain in the limit $t\rightarrow 0^{+}$, we must consider this exponential contribution and observe where it switches on across the real-$x$ axis. To demonstrate how this works, we follow the framework in [20], summarise the main results here, and provide further details in appendix A (see [21, 67, 68, 69, 91], for example, for other studies of wave motion using this framework of exponential asymptotics).
A crucial step is to analyse the late terms in (2.4), which satisfy
$(n+1)u_{n+1}=-6\sum_{j=0}^{n}u_{j}u^{\prime}_{n-j}-u_{n}^{\prime\prime\prime}.$ (2.5)
Due to the factorial over power divergence of (2.4), following Dingle [37] we apply the ansatz
$u_{n}\sim\frac{\mathcal{A}(x)\Gamma(2n+\gamma)}{\chi(x)^{2n+\gamma}}\quad\mbox{as}\quad n\rightarrow\infty,$ (2.6)
which, after substituting into (2.5), leads to
$\frac{1}{2}\chi=\left(\chi^{\prime}\right)^{3},\quad\frac{\gamma}{2}\mathcal{A}=3\mathcal{A}\chi^{\prime}\chi^{\prime\prime}+3\mathcal{A}^{\prime}\left(\chi^{\prime}\right)^{2}.$ (2.7)
Here $\chi$ is the so-called singulant, which must vanish at singularities of $u_{0}$, and it suffices to consider the location $x=x_{0}$ (the full expansion of $u_{n}$ for large $n$ comprises a sum of terms of the form (2.6) associated with each singularity $x_{0}$). Thus, given $x_{0}$ is assumed to lie in the upper half plane, we explain in appendix A.1 that the appropriate solutions are
$\chi=-\frac{2}{3^{3/2}}(x-x_{0})^{3/2},\quad\mathcal{A}=\Lambda(x-x_{0})^{(\gamma-1)/2},$ (2.8)
@@
$u_{n}\sim\frac{\Lambda(-3^{3/2}/2)^{2n+\gamma}\Gamma(2n+\gamma)}{(x-x_{0})^{3n+\gamma+1/2}}\quad\mbox{as}\quad n\rightarrow\infty.$ (2.9)
For the initial conditions we are concerned with in this study, we have (1.3). To be consistent with this local behaviour near $x=x_{0}$, we must choose $\gamma=3/2$.
@@
$\pi\mathrm{i}\,\mathcal{A}\,t^{-\gamma/2}\mathrm{e}^{-\chi/t^{1/2}}=\pi\mathrm{i}\,\Lambda(x-x_{0})^{1/4}t^{-3/4}\mathrm{e}^{2(x-x_{0})^{3/2}/3(3t)^{1/2}}$ (2.10)
switches on as we cross the Stokes line $\mathrm{arg}(x-x_{0})=-2\pi/3$ from right to left (given that $x_{0}$ is in the upper half $x$ plane). This Stokes line is found by setting the singulant term $\chi$ to be real and positive. There will be another term like (2.10) switched on across a Stokes line born in the lower half plane, and together the sum of these provides the asymptotic behaviour of the dispersive waves on the real line in the small-time limit.
To determine the constant $\Lambda$ in (2.10), we need to match into an inner region near $x=x_{0}$, which we study below in subsection 3.1. The details for this matching are provided in appendix A.3, with the key result that
$\Lambda=\frac{\mathrm{i}}{3^{3/4}\pi^{3/2}}\cos\left(\frac{\pi}{2}\sqrt{1-4A_{0}}\right).$ (2.11)
A striking property of $\Lambda$ is that it vanishes for $A_{0}=-N(N+1)$, where $N\in\mathbb{N}$. Therefore, this leading-order result predicts that a train of dispersive waves is associated with each complex conjugate pair of singularities of the initial condition $u_{0}$, with the amplitude vanishing for $A_{0}=-N(N+1)$. This prediction is consistent with the $\mathrm{sech}^{2}$-type initial condition (1.4), for which it is known that there are no dispersive waves when $A_{0}$ takes on these triangular values (these are the $N$-soliton solutions, having flat tails). (More generally, if $A_{0}=-N(N+1)$ for an initial condition that is not of the special form (1.4), then parts of the required asymptotics to describe the dispersive waves will be different; we include a discussion on such cases in appendix A.4.)
To take an example, consider an initial condition that has a singularity lying on the imaginary axis, which we write as $x_{0}=\mathrm{i}y_{0}$. Our analysis predicts that the term (2.10) switches on across the Stokes line $\mathrm{arg}(x-\mathrm{i}y_{0})=-2\pi/3$ in the upper half $x$ plane. We would also observe an analogous term that switches on across $\mathrm{arg}(x+\mathrm{i}y_{0})=2\pi/3$ in the lower half plane. These Stokes lines both hit the real axis at $x=-y_{0}/\sqrt{3}$. Thus, putting it together, we find the exponential contribution to our asymptotic solution on the real-$x$ axis is of the form